OpenStudy (anonymous):

C3H8 + 5O2 → 3CO2 + 4H2O How many moles of H2O are formed if 5 moles of propane react with oxygen gas?

OpenStudy (somy):

ratio game look at coefficients before propane and H2O compounds its 1 for propane and 4 for H2O right? so this is the ratio 1:4 so lets see 1 propane is 5 mole now how much is 4? 1-5 4- X X= 4*5 = .....mole of H2O @aaronq am i getting this questing right?

OpenStudy (anonymous):

yes thank you

OpenStudy (aaronq):

The easiest way to see it if you make fractions \(\sf \dfrac{moles~of~propane}{1}=\dfrac{moles~of~water}{4}\) \(\sf \dfrac{5}{1}=\dfrac{moles~of~water}{4}\) moles of water = \(\sf \dfrac{5*4}{1}\)

OpenStudy (anonymous):

4Fe + 3O2 → 2Fe2O3, if 8 moles of iron and excess oxygen from the air were available, how many moles of iron (III) oxide would be produced? need help ?

OpenStudy (anonymous):

@aaronq

OpenStudy (somy):

well its same thing lol

OpenStudy (aaronq):

^it is the same thing, make the fractions/ratios first

OpenStudy (anonymous):

thank you